OpenStudy (anonymous):

If x and y vary directly and x = 5 when y = -15, what is the value of x when y = 3?

OpenStudy (anonymous):

vary means a ration of x/y or y/x. 5/-15 = -(1/3) if y = 3 x = -(1/3) * 3 x = -1